What is the distance to a star if the parallax is 0.022 arc-seconds? Give your answer to 1 decimal place


Note: Answer was 45.5, but I just don't get it
Answer #1

See that we need to calculate distance to a star from given parallax

Now equation is given by:

d = 1/p

d = distance to a star (in parsec)

p = parallax = 0.022 arc-seconds

So Using this value

d = 1/p = 1/0.022 = 45.4545

Since we need answer in one decimal place, So

d = 45.5 parsec

Please Upvote.

Comment below if you have any query.